# LaViPlan : Language-Guided Visual Path Planning with RLVR

### v2

@@ 4.2.2 Effective Sample Set for GRPO
Table 5: Performance impact of different easy-to-hard data ratios when the total number of training samples is held constant.
@@
9:1 3.84 (-15.0%) 5.05 (+10.7%) 4.09 (-8.3%) 6.31 (-2.3%)
7:3 5.55 (+22.8%) 6.70 (+18.4%) 4.05 (-9.2%) 6.16 (-4.6%)
6:4 3.62 (-19.9%) 4.83 (-14.7%) 3.85 (-13.7%) 6.09 (-5.7%)
To evaluate the impact of sample difficulty on reinforcement fine-tuning, we fix the SFT:RFT ratio at 4:1 (corresponding to 4,344 and 1,086 samples, respectively), and vary the proportion of easy and hard samples within this fixed budget. As shown in Tab. 5, allocating a larger portion of hard scenarios (up to 40%) leads to consistent improvements across both ADE and FDE, particularly for hard cases. These results suggest that the reinforcement fine-tuning benefits from challenging samples that provide richer learning signals, helping the model generalize better in complex planning situations.
@@ 4.2.3 Impact of Reasoning in Two-Phased Fine-tuning
Baseline 4.52 5.66 4.46 6.46
w/o reasoning 3.71 (-17.7%) 4.95 (-12.6%) 3.96 (-11.2%) 6.26 (-3.1%)
w/ reasoning 3.62 (-19.9%) 4.83 (-14.7%) 3.85 (-13.7%) 6.09 (-5.7%)
5 Discussion and Limitations
We proposed LaViPlan, a reinforcement learning framework guided by planning-oriented metrics, as a method for fine-tuning VLMs on autonomous driving tasks. Our findings suggest that RLVR not only enhances the zero-shot scene understanding capabilities of LLM agents but also shows potential in mitigating the misalignment between vision, language, and action. In conclusion, several limitations remain and highlight directions for future research. (1) The planning reward is sparse, as it is only obtained after the entire rollout, which highlights the need for step-wise feedback during policy optimization. This may have constrained the potential of GRPO, limiting the effectiveness of intermediate decision-making signals [32]. (2) While the improvements in planning performance indicate progress, they are not sufficient to claim complete resolution of vision-language-action misalignment. A more context-aware evaluation metric is needed to assess whether model decisions are grounded with context-awareness [29]. (3) The extent to which this effectiveness transfers to world models remains an open question—particularly in the realm of counterfactual reasoning over out-of-distribution (OOD) actions encountered in sequential decision-making tasks.
6 Acknowledgement
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2023-00236245, Development of Perception/Planning AI SW for Seamless Autonomous Driving in Adverse Weather/Unstructured Environment.
Appendix
@@ B.1 Dataset Construction and Splitting Strategy
  • Top-$N$ samples with high variance $\rightarrow$ turning samples
@@
To enable both supervised and reinforcement fine-tuning for visual path planning, we split the dataset into Supervised Fine-Tuning (SFT) and Reinforcement Fine-Tuning (RFT) subsets using a structured and variance-aware strategy. We first load the complete set of trajectory metadata $D$ and determine the target ratios for SFT versus RFT (e.g., 4:1), as well as the distribution of turning and straight trajectories within each subset (e.g., 6:4 or 4:6 depending on the setting). For each trajectory $d\in D$, we calculate the variance of its $x$-coordinates, which serves as a proxy for trajectory curvature. We then sort the trajectories in descending order of $x$-variance. Based on this ordering, we classify the top-$N$ trajectories with high variance as turning trajectories and the remaining as straight trajectories. From these, we allocate subsets to SFT and RFT splits: a portion of turning trajectories to RFT-turn and the remainder to SFT-turn, and similarly, straight trajectories to SFT-straight and RFT-straight. Each selected sample is then converted into an instruction-following format consisting of an image reference, a natural language prompt, and a corresponding reasoning-answer pair. These formatted samples are saved as separate JSON files for the SFT and RFT datasets.
@@ B.2 Validation Set Construction
For evaluation, we construct validation sets focused on moderate and hard scenarios. We begin with a dense set of validation trajectories $D_{\text{dense}}$ about 12K samples and remove any overlap with standard annotations $D_{\text{standard}}$ to obtain candidate validation samples $D_{\text{val}}$. We compute the $x$-variance for each sample and sort them accordingly. To form the easy set $D_{\text{easy}}$, we select a middle slice (e.g., 1K samples centered around the median $x$-variance). For the hard set $D_{\text{hard}}$, we randomly sample 0.7K trajectories from the top 70% (high-variance) and 0.3K from the bottom 10% (low-variance) of the sorted list. Both validation subsets are saved as JSON files for downstream evaluation.
Algorithm 2 Construct Validation Sets (Easy, Hard)
1:Load dense validation trajectories $D_{dense}$
2:Load standard validation annotations $D_{standard}$
3:$D_{val}\leftarrow D_{dense}\setminus D_{standard}$
4:For each $d\in D_{val}$, compute variance of $x$ over trajectory
5:$D_{sorted}\leftarrow$ sort $D_{val}$ by $x$-variance (descending)
6:$N\leftarrow|D_{sorted}|$
7:Easy Set: $D_{easy}\leftarrow D_{sorted}[N/2-500:N/2+500]$
8:Hard Set:
9:Randomly sample 700 from top 70% of $D_{sorted}$
10:Randomly sample 300 from bottom 10% of $D_{sorted}$
11:$D_{hard}\leftarrow$ combined samples
12:Save $D_{easy}$ and $D_{hard}$ as JSON

### v4

@@ 4.3.2 Effective Sample Set for GRPO
To evaluate the impact of sample difficulty on reinforcement fine-tuning, we fix the SFT:RFT ratio at 4:1 (corresponding to 4,344 and 1,086 samples, respectively), and vary the proportion of easy and hard samples within this fixed budget.
In-Domain Dataset. As shown in Tab. 6, allocating a larger portion of hard scenarios (up to 40%) leads to consistent improvements across both ADE and FDE, particularly for hard cases. These results suggest that reinforcement fine-tuning benefits from challenging samples that provide richer learning signals, thereby enhancing the model’s generalization ability in complex planning situations. Notably, the 6:4 ratio achieves the best overall performance, as it strikes an effective balance between stable learning from easy cases and the additional supervision obtained from hard cases. While increasing the hard portion further (e.g., 7:3) introduces more diverse challenges, it can undermine the stability of learning in easy cases, resulting in less consistent improvements across metrics.
Table 6: Performance impact of different easy-to-hard data ratio when the total number of training samples is held constant with in-domain dataset.
@@
9:1 3.84 (-6.8%) 5.05 (-4.9%) 4.09 (-7.9%) 6.31 (-3.1%)
7:3 5.55 (+34.7%) 6.70 (+26.2%) 4.05 (-8.8%) 6.16 (-5.4%)
6:4 3.62 (-12.1%) 4.83 (-9.1%) 3.85 (-13.3%) 6.09 (-6.5%)
Out-of-Domain Dataset. The ground-truth trajectory does not always align with the safest path, and such misalignment becomes more prominent in difficult scenarios. In this setting, the 7:3 ratio yields the most favorable outcomes as shown in Tab. 7, especially under safety-focused and balanced metrics. This is because the larger share of hard samples exposes the model to safety-critical conflicts, encouraging it to learn beyond simply imitating labeled trajectories and to acquire strategies that prioritize safe driving. However, further increasing the proportion of hard samples to 6:4 biases the training distribution excessively toward complex cases, which diminishes generalization in performance-focused metrics.
Table 7: Performance impact of different easy-to-hard data ratio with out-of-domain dataset and absolute difference from baseline.
Ratio / Model Balanced $\uparrow$ Safety-Focused $\uparrow$ Performance-Focused $\uparrow$ Equal $\uparrow$
SFT (5K) 0.60 (+0.20) 0.59 (+0.29) 0.56 (+0.06) 0.63 (+0.30)
LaViPlan (9:1) 0.58 (+0.18) 0.62 (+0.32) 0.51 (+0.01) 0.63 (+0.30)
LaViPlan (7:3) 0.64 (+0.24) 0.73 (+0.43) 0.56 (+0.06) 0.70 (+0.37)
LaViPlan (6:4) 0.45 (+0.05) 0.49 (+0.19) 0.39 (-0.11) 0.51 (+0.18)
In short, our findings indicate that the effectiveness of easy-to-hard sampling is highly domain-sensitive, suggesting that in-domain performance benefits from a balanced exposure to both easy and hard cases, while out-of-domain robustness requires greater emphasis on hard cases to capture safety-critical cues.
@@ 4.3.3 Impact of Reasoning in Two-Phased Fine-tuning
Baseline 4.12 5.31 4.44 6.51
w/o reasoning 3.71 (-9.95%) 4.95 (-6.78%) 3.96 (-10.81%) 6.26 (-3.84%)
w/ reasoning 3.62 (-12.14%) 4.83 (-9.04%) 3.85 (-13.29%) 6.09 (-6.45%)
5 Discussion
Sparse Reward. RLVR can guarantee performance improvement, but its efficiency may suffer when rewards are sparse. In our setting, the planning reward is sparse, as ADE is only computed after the entire rollout, emphasizing the importance of step-wise feedback during policy optimization. This sparsity may have limited GRPO’s potential, reducing the effectiveness of intermediate decision-making signals [37]. Moreover, sparse rewards can slow convergence and increase variance in policy updates, suggesting that integrating denser or auxiliary feedback could further improve learning efficiency and the quality of generated trajectories [24].
GRPO as Policy Regularizer, RLVR as Affordance. GRPO-based fine-tuning is not a silver bullet and relies on the presence of a reasonably strong SFT model as a prerequisite. This is because the policy model is optimized under a divergence constraint that keeps it close to the baseline model. Moreover, the reward function designed for reinforcement fine-tuning defines the affordance of the model, shaping which behaviors are reinforced. Therefore, beyond incorporating existing metrics into the reward design, future work should consider reward formulations that explicitly account for safety, success rate, and other critical factors, rather than solely imitating the ground-truth trajectory.
6 Conclusion
In this work, we present LaViPlan, a reinforcement learning framework guided by planning-oriented metrics for fine-tuning VLMs on autonomous driving tasks. Our findings indicate that RLVR enhances the zero-shot scene understanding capabilities of VLMs and may help mitigate misalignment between vision, language, and action. However, the extent to which these benefits generalize to world models—particularly in counterfactual reasoning under actions outside the training distribution in sequential decision-making tasks—remains an open question.
7 Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2023-00236245, Development of Perception/Planning AI SW for Seamless Autonomous Driving in Adverse Weather/Unstructured Environment)
@@ B.1 Dataset Construction and Splitting Strategy
  • Top-NN samples with high variance $\rightarrow$ turning samples
@@
To enable both supervised and reinforcement fine-tuning, we split the dataset into Supervised Fine-Tuning (SFT) and Reinforcement Fine-Tuning (RFT) subsets using a structured and variance-aware strategy. We first load the complete set of trajectory metadata $D$ and determine the target ratios for SFT versus RFT (e.g., 4:1), as well as the distribution of turning and straight trajectories within each subset (e.g., 6:4 or 4:6 depending on the setting). For each trajectory $d\in D$, we calculate the variance of its $x$-coordinates, which serves as a proxy for trajectory curvature. We then sort the trajectories in descending order of $x$-variance. Based on this, we classify the top-$N$ trajectories with high variance as turning trajectories and the remaining as straight trajectories. From these, we allocate subsets to SFT and RFT splits: a portion of turning trajectories to RFT-turn and the remainder to SFT-turn, and similarly, straight trajectories to SFT-straight and RFT-straight. Each selected sample is then converted into an instruction-following format consisting of an image reference, a natural language prompt, and a corresponding reasoning-answer pair.
@@ B.2 Validation Set Construction
For evaluation, we construct validation sets focused on moderate and hard scenarios as shown in Tab. 11. We begin with a dense set of validation trajectories $D_{\text{dense}}$ about 12K samples and remove any overlap with standard annotations $D_{\text{standard}}$ to obtain candidate validation samples $D_{\text{val}}$. We compute the $x$-variance for each sample and sort them accordingly. To form the easy set $D_{\text{easy}}$, we select a middle slice (e.g., 1K samples centered around the median $x$-variance). For the hard set $D_{\text{hard}}$, we randomly sample 0.7K trajectories from the top 70% (high-variance) and 0.3K from the bottom 10% (low-variance) of the sorted list.
Algorithm 2 Construct Validation Sets (Easy, Hard)
1:Load dense validation trajectories $D_{dense}$
2:Load standard validation annotations $D_{standard}$
[…]
4:For each $d\in D_{val}$, compute variance of $x$ over trajectory
[…]
7:Easy Set: $D_{easy}\leftarrow D_{sorted}[N/2-500:N/2+500]$
8:Hard Set:
9:Randomly sample 700 from top 70% of $D_{sorted}$
10:Randomly sample 300 from bottom 10% of $D_{sorted}$
[…]
12:Save $D_{easy}$ and $D_{hard}$ as JSON
@@
Appendix C Pseudo code of OOD Evaluation
Algorithm 3 Procedure for Out-of-Distribution evaluation with as pseudo code.
1:Dataset $\mathcal{D}=\{(I_{i},B_{i})\}_{i=1}^{N}$ where $I_{i}$: image, $B_{i}$: 2D bounding boxes
2:Model $\mathcal{M}$ predicting trajectory $\mathcal{T}=\{(x_{j},y_{j})\}_{j=1}^{20}$
3:Fail Rate, Avg 2D BBox Collision, Avg Penetration Length
4:Initialize: $C_{fail}\leftarrow 0$, $C_{bbox}\leftarrow 0$, $L_{pen}\leftarrow 0$, $N\leftarrow 0$, $T_{bbox}\leftarrow 0$
5:for each $(I,B)$ in $\mathcal{D}$ do
6:  $\mathcal{T}\leftarrow\mathcal{M}(I)$ $\triangleright$ 20-point trajectory
7:  $traj\leftarrow$ LineString($\mathcal{T}$), $c\leftarrow 0$
8:  for each $b\in B$ do
9:   $p\leftarrow$ Polygon($b$)
10:   if trajtraj.intersects(ppthen
11:     $c\leftarrow c+1$
12:     $L_{pen}\leftarrow L_{pen}+\text{length}(traj\cap p)$
13:   end if
14:   $T_{bbox}\leftarrow T_{bbox}+1$
15:  end for
16:  if $c>0$ then $C_{fail}\leftarrow C_{fail}+1$
17:  end if
18:  $C_{bbox}\leftarrow C_{bbox}+c$, $N\leftarrow N+1$
19:end for
[…]
Trajectory Fail Rate: The proportion of predicted trajectories that intersect with at least one 2D bounding box in the scene:
$\text{Fail Rate}=\frac{\text{Number of collision trajectories}}{\text{Total number of samples}}$ (8)
Average 2D BBox Collision: The average number of bounding boxes that each predicted trajectory collides with:
$\text{Collision Count}=\frac{\text{Total bbox collisions}}{\text{Total number of samples}}$ (9)
Penetration Length: The average geometric length of trajectory segments that penetrate into bounding boxes:
$\text{Avg Penetration Length}=\frac{\sum_{i=1}^{N}\sum_{j=1}^{|B_{i}|}\text{length}(\mathcal{T}_{i}\cap bbox_{j})}{\text{Total number of samples}}$ (10)
where $\mathcal{T}_{i}\cap bbox_{j}$ represents the intersection between trajectory $\mathcal{T}_{i}$ and bounding box $bbox_{j}$.
Appendix D Qualitative Results on OOD Benchmark
We qualitatively evaluate the proposed method on CODA-LM, a zero-shot dataset involving diverse corner cases and road hazards. Since CODA-LM does not provide ground-truth trajectories, we rely on visual inspection of the predicted paths to assess plausibility, safety, and contextual awareness. As shown in Fig. 6, the model fine-tuned with verifiable rewards produces more realistic and contextually appropriate trajectories compared to the one trained via supervised fine-tuning. In scenarios involving roadwork, blocked lanes, and ambiguous path constraints, the supervised model often generates linear or risk-prone trajectories that lack appropriate deviation or caution. In contrast, the proposed method consistently generates smoother, safer, and more context-aware trajectories, often adjusting path curvature to avoid obstacles such as cones, barriers, and vehicles. These results demonstrate the model’s capacity to generalize to previously unseen situations by aligning its reasoning and path generation with latent planning cues in the scene, even without explicit supervision. This qualitative result suggests that reinforcement fine-tuning enhances the model’s ability to adapt to driving contexts.